Delph Company uses a job-order costing system and has two manufacturing departments-Molding and Fabrication. The company provided the following estimates at the beginning of the year: Tota 70,800 Molding Fabrication Machine-hours Fixed manufacturing overhead costs Variable manufacturing overhead cost per machine-hour 49,000 730,000 300,000 $1,030,900 30,000 $5.60 5.6 During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs- Job D-70 and Job C-200. It provided the following information related to those two jobs: Job D-70: Direct materials cost Direct labor cost Machine-hours Molding Fabrication Total S376,000 325,000 701,0e0 21e,e00 140,00 350,9e0 22,800 8,000 30,000 Job C-200: Direct materials cost Direct labor cost Machine-hours Molding Fabrication Total 300,000 220,000 520,90e 140,00e 280,000 420,00e 8,000 32,000 40,000 Delph had no underapplied or overapplied manufacturing overhead during the year Exercise 2-15 Part 2 2. Assume Delph uses departmental predetermined overhead rates based on machine-hours. a. Compute the departmental predetermined overhead rates b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200 C. lf Delph establishes bid prices that are 120% of total manufacturing costs, what bid prices would it have established for Job D-70 and Job C-200? d. What is Delphs cost of goods sold for the year?
Answer #1

a) Department overhead rate :

Molding department = (730000/30000)+5.60 = $29.93 per machine hour

Fabrication department = (300000/40000)+5.60 = $13.10 per machine hour

b) Calculate total manufacturing cost

Job D-70 Job C-200
Direct material 701000 520000
Direct labor 350000 420000
Overhead applied
Molding 22000*29.93 = 658533 8000*29.93 = 239467
Fabrication 8000*13.10 = 104800 32000*13.10 = 419200
Total manufacturing cost 1814333 1598667

c) Bid price

Job D-70 = 1814333*1.2 = $2177200

Job C-200 = 1598667*1.2 = $1918400

d) Cost of goods sold = 1814333+1598667 = $3413000